OpenStudy (anonymous):

The Everton College store paid $1587 for an order of 45 calculator. The store paid $11 for each scientific calculator. The other, all graphin calculators, cost the store $53 each. How many of each type of calculator was order?

OpenStudy (hoblos):

let x= number of scientific calculator y= number of graphing calculators x+y = 45 -----> x=45-y 11x+53y = 1587 11(45-y) +53y = 1587 495 -11y +53y = 1587 42y = 1092 y = 26 x=45-y = 45-26 = 19

OpenStudy (anonymous):

thanks

OpenStudy (hoblos):

my pleasure XD